# On a Super-analogue of Carrollian Manifolds

### v2

@@ 2.3. The Super-Carrollian Lie Algebra
Let $(M,g,Q,P)$ be a super-Carrollian manifold. Then $Q$ cannot be a Killing vector field, i.e., $\mathcal{L}_{Q}g\neq 0$.
@@
$(L_{Q}g)(Q,X)=Q\langle Q|X\rangle-\langle[Q,Q]|X\rangle+\langle Q|[Q,X]\rangle=-2\,\langle P|X\rangle\,.$
Thus, as $P$ is not in the kernel of $g$ we cannot have $(L_{Q}g)(Q,X)=0$ for all $X$, and so $L_{Q}g\neq 0$. ∎
@@
Let $(M,g,Q,P)$ be a super-Carrollian manifold. The odd vector field $Q$ is not an element of $\mathfrak{scarr}_{1}(M,g,Q,P)$.
@@
We can interpret Proposition 2.21 as saying that if $Q$ were killing, then $\ker(g)$ must include $P$; and so the rank of the kernel must be (at least) $1|1$. However, this directly violates the requirements of the definition of a super-Carrollian manifold. It is important to note that $P$ may be Killing, and this does not destroy the super-Carrollian manifold structure. If $P$ is a Killing vector field, then the local components of the degenerate metric in Shander coordinates are independent of the even coordinate $t$.
@@
Let $(M,g,Q,P)$ be a super-Carrollian manifold and let $\mathfrak{scarr}(M,g,Q,P)$ be its super-Carrollian Lie algebra.
@@
    The Lie algebra $\mathfrak{scarr}_{0}(M,g,Q,P)$ is isomorphic to the Lie algebra of Killing vector fields of $(M_{red},g_{red})$ that commute with $P_{red}$.
@@
    If $P$ is Killing, then $\mathfrak{scarr}_{1}(M,g,Q,P)=\operatorname{Span}_{{\mathbb{R}}}\{D\}$, otherwise $\mathfrak{scarr}_{1}(M,g,Q,P)=\{\mathbf{0}\}$.
@@
    • If $P$ is Killing, then $\partial_{t}g_{ba}=0$, $\partial_{t}g_{ta}=0$, and $\partial_{t}g_{tt}=0$. Then for $Y$ to be Killing we require $\partial_{a}f=0$, and so $f=\mathrm{const}$. Thus, $\mathfrak{scarr}_{1}(M,g,Q,P)=\operatorname{Span}_{{\mathbb{R}}}\{D\}$.
    • If $P$ is not Killing, then $\partial_{t}g_{ba}\neq 0$, $\partial_{t}g_{ta}\neq 0$, and $\partial_{t}g_{tt}\neq 0$. Then for $Y$ to be Killing the condition $f=0$ is forced. Thus, $\mathfrak{scarr}_{1}(M,g,Q,P)=\{\mathbf{0}\}$.
Provided $n\neq 2$ (see Proposition 2.18), the dimension of the Lie algebra of $g_{red}$ is bounded by $n(n+1)/2$, we have the following corollary of Proposition 2.25.
@@
Let $(M,g,Q,P)$ be a super-Carrollian manifold with $\dim M_{red}\neq 2$, then the Lie superalgebra $\mathfrak{scarr}(M,g,Q,P)$ is finite-dimensional.
@@
The finite dimensionality of $\mathfrak{scarr}(M,g,Q,P)$ is to be contrasted with infinitesimal automorphisms of weak Carrollian manifolds where the Lie algebra is infinite-dimensional (see Duval et al. [10]).
@@
A super-Carrollian manifold $(M,g,Q,P)$ such that $P$ is Killing, i.e., $\mathcal{L}_{P}g=0$, is referred to as a static super-Carrollian manifold.
Note that for any static super-Carrollian manifold $P\in\mathfrak{scarr}_{0}(M,g,Q,P)$ as $\mathcal{L}_{P}Q=[P,Q]=0$. Thus, the super-Carrollian Lie algebra of a static super-Carrollian manifold consists of at least $P$ and $D$. As standard, we have a representation of the $N=1$, $d=1$ supertranslation algebra given by [D,D]=2P[D,D]=-2P and $[P,Q]=0$. Thus, we have established the following.
@@
Let $(M,g,Q,P)$ be a static super-Carrollian manifold, then the $N=1$, $d=1$ supertranslation algebra is a Lie subsuperalgebra of $\mathfrak{scarr}_{(}M,g,Q,P)$ realised by the vector fields $P$ and $D$.
@@
i.e., we have the usual Minkowski spacetime of signature $(3,1)$. Clearly, $P$ is Killing and we have a static super-Carrollian manifold, so the super-Carrollian Lie algebra contains $D$. The isometries of $g_{red}$ are given by the Poincaré Lie algebra, i.e., $\mathfrak{iso}(g_{red})\simeq\mathfrak{so}(3,1)\ltimes{\mathbb{R}}^{4}$. However, we require the Lie subalgebra generated by $P_{red}=\partial_{t}$ and the other generators of the Poincaré Lie algebra that commute with $P_{red}$. The remaining transformations are spacial rotations, spacial translations, and temporal translations. Importantly, there are no boosts. Thus,
@@
Let $(M_{0},g_{0})$ be a (pseudo-)Riemannian manifold with $\dim M_{0}\geq 2$, and let $\mathfrak{iso}(g_{0})$ be its isometry Lie algebra. We then equip $M_{red}:=M_{0}\times{\mathbb{R}}$ with the product metric $g_{red}:=g_{0}\oplus g_{\mathbb{R}}$, where $g_{\mathbb{R}}$ is the standard constant metric on $\mathbb{R}$}. An established result is that $\mathfrak{iso}(g_{0}\oplus g_{\mathbb{R}})\simeq\mathfrak{iso}(g_{0})\oplus\mathfrak{iso}(g_{\mathbb{R}})$. Recall that $\mathfrak{iso}(g_{R})$ consists of just translations and thus is identified with $\mathfrak{u}(1)$. Then $M:=M_{0}\times{\mathbb{R}}^{1|1}$ is a super-Carrollian manifold whose degenerate metric written in Shander coordinates is
@@
$\mathfrak{scarr}(M_{0}\times{\mathbb{R}}^{1|1})\simeq\big{(}\mathfrak{iso}(M_{0})\oplus\mathfrak{u}(1)\big{)}\oplus_{\textrm{E}xt}{\mathbb{R}}^{0|1}\,$,
@@
2.4. Connections on Super-Carrollian Manifolds
@@
Let $(M,g,Q,P)$ be a super-Carrollian manifold. An affine connection on the supermanifold $M$ is said to be
@@
Let $(M,g,Q,P)$ be a super-Carrollian manifold equipped with an affine connection $\nabla$. If the affine connection is supersymmetry compatible, then $\nabla$ cannot be torsionless.
@@
Let $(M,g,Q,P)$ be a super-Carrollian manifold equipped with an affine connection $\nabla$. If the affine connection is supersymmetry compatible, then $Q\in\ker\big{(}R_{\nabla}(X,Y)\big{)}$ for all pairs of vector fields $X,Y\in{Vect}(M)$.
@@
Supersymmetry compatible connections exist on any super-Carrollian manifold.
@@
Let $(M,g,Q,P)$ be a super-Carrollian manifold equipped with a metric compatible connection. For every $X\in{Vect}(M)$ there exists a function $f_{X}\in C^{\infty}(M)$ ($\widetilde{f_{X}}=\widetilde{X}$), such that $\nabla_{X}Q=f_{X}\,Q$.
@@
Let $(M,g,Q,P)$ be a super-Carrollian manifold equipped with an affine connection $\nabla$. If the affine connection is metric compatible, then $\nabla$ cannot be torsionless.
@@
Let $(M,g,Q,P)$ be a super-Carrollian manifold equipped with an affine connection. If the affine connection is metric compatible, then $Q$ is a generalised eigenvector of $R_{\nabla}(X,Y)$ for all pairs of vector fields $X,Y\in{Vect}(M)$.
@@
does not fully determine a metric compatible connection. While $\langle\nabla_{X}Y|Z\rangle$ is fully determined for a given metric compatible connection (assuming at least one exists), $\nabla_{X}Y$ is only determined up to a vector in $\ker(g)$. Thus, we cannot expect a direct analogue of the fundamental theorem of Riemannian supergeometry for super-Carrollian manifolds. Nonetheless, we have the following theorem.
@@
Metric compatible affine connections exist on any super-Carrollian manifold.
@@
Compatible affine connections exist on any super-Carrollian manifold.
@@ 3. Concluding Remarks
An odd analogue of a Carrollian manifold has been constructed and studied. We have shown that supersymmetry compatible and metric compatible connections exist on any super-Carrollian manifold, and importantly, that compatible connections always exist, i.e., affine connections that are both supersymmetry compatible and metric compatible can always be constructed. It is the freedom in defining a connection, thanks to the degeneracy of the metric, that allows these two conditions to be simultaneously satisfied. It was argued by Bekaert & Morand [4] that only invariant Carrollian manifolds, i.e., Carrollian manifolds for which the fundamental/Carrollian vector field is Killing, can admit torsion-free compatible connections. This is in stark contrast with super-Carrollian manifolds, where the supersymmetry generator $Q$ cannot be Killing, and compatible connections must carry torsion. The two geometries are fundamentally different; the non-integrability of the kernel of the super-case is the root of these differences.
The Lie superalgebra of infinitesimal automorphisms of a super-Carrollian manifold, which we referred to as the super-Carrollian Lie algebra, has been studied. An interesting result is that this Lie superalgebra is finite-dimensional and tightly tied to the Lie algebra of Killing vector fields of the reduced metric (modulo a complication in two even dimensions). In the classical setting of weak Carrollian manifolds, the Lie algebra of infinitesimal automorphisms is infinite-dimensional. When the extra condition of preserving a compatible connection is imposed, the Lie algebra becomes finite-dimensional; see Duval et al. [10] and references therein for details.
While the main motivation for this work is rooted in mathematical curiosity, further explicit examples of super-Carrollian manifolds are desirable and could expose applications thereof in physics. It may be possible to formulate a superspace version of non-holonomic supermechanics where the dynamics generated by $Q$ could depend on the extra even coordinates understood as “external parameters”; which might be interpreted as the classical background fields or slowly varying time-dependent parameters of the physical system. This could provide a novel way to study supersymmetric mechanical systems with evolving external parameters. For example, there may be supersymmetric analogues of geometric phases, such as the Hannay angle (see [18]), underlying these models. Another direction of investigation is to find strong links between super-Carrollian manifolds and super-Carrollian field theories, see [3, 20, 25, 33], for example. There is also the possibility of applications in condensed matter physics, where, for example, Carrollian physics appears in magic bi-layer graphene (see [1]). More speculatively, super-Carrollian manifolds could provide the geometric framework for understanding tensionless superstrings and their Carrollian symmetries, asymptotic supersymmetries, super-BMS symmetries, and Carrollian supergravity and holography.

### v3

@@ 2.3. The Super-Null Lie Algebra
Let $(M,g,Q,P)$ be a super-null Riemannian manifold. Then $Q$ cannot be a Killing vector field, i.e., $\mathcal{L}_{Q}g\neq 0$.
@@
$(\mathcal{L}_{Q}g)(Q,X)=Q\langle Q|X\rangle-\langle[Q,Q]|X\rangle+\langle Q|[Q,X]\rangle=-2\,\langle P|X\rangle\,.$
Thus, as $P$ is not in the kernel of $g$ we cannot have $(\mathcal{L}_{Q}g)(Q,X)=0$ for all $X$, and so $\mathcal{L}_{Q}g\neq 0$. ∎
@@
Let $(M,g,Q,P)$ be a super-null Riemannian manifold. The odd vector field $Q$ is not an element of $\mathfrak{snull}_{1}(M,g,Q,P)$.
@@
We can interpret Proposition 2.4 as saying that if $Q$ were killing, then $\ker(g)$ must include $P$; and so the rank of the kernel must be (at least) $1|1$. However, this directly violates the requirements of the definition of a super-null Riemannian manifold. It is important to note that $P$ may be Killing, and this does not destroy the super-null Riemannian manifold structure. If $P$ is a Killing vector field, then the local components of the degenerate metric in Shander coordinates are independent of the even coordinate $t$.
@@
Let $(M,g,Q,P)$ be a super-null Riemannian manifold and let $\mathfrak{snull}(M,g,Q,P)$ be its super-null Lie algebra.
@@
    The Lie algebra $\mathfrak{snull}_{0}(M,g,Q,P)$ is isomorphic to the Lie algebra of Killing vector fields of $(M_{red},g_{red})$ that commute with $P_{red}$.
@@
    If $P$ is Killing, then $\mathfrak{snull}_{1}(M,g,Q,P)=\operatorname{Span}_{{\mathbb{R}}}\{D\}$, otherwise $\mathfrak{snull}_{1}(M,g,Q,P)=\{\mathbf{0}\}$.
@@
    • If $P$ is Killing, then $\partial_{t}g_{ba}=0$, $\partial_{t}g_{ta}=0$, and $\partial_{t}g_{tt}=0$. Then for $Y$ to be Killing we require $\partial_{a}f=0$, and so $f=\mathrm{const}$. Thus, $\mathfrak{snull}_{1}(M,g,Q,P)=\operatorname{Span}_{{\mathbb{R}}}\{D\}$.
    • If $P$ is not Killing, then $\partial_{t}g_{ba}\neq 0$, $\partial_{t}g_{ta}\neq 0$, and $\partial_{t}g_{tt}\neq 0$. Then for $Y$ to be Killing the condition $f=0$ is forced. Thus, $\mathfrak{snull}_{1}(M,g,Q,P)=\{\mathbf{0}\}$.
As $(M_{red},g_{red})$ is a pseudo-Riemannian manifold (see Proposition 2.3), the dimension of the Lie algebra of $g_{red}$ is bounded by $n(n+1)/2$, we have the following corollary of Proposition 2.5.
@@
Let $(M,g,Q,P)$ be a super-null Riemannian manifold, then the Lie superalgebra $\mathfrak{snull}(M,g,Q,P)$ is finite-dimensional.
@@
The finite dimensionality of $\mathfrak{snull}(M,g,Q,P)$ is to be contrasted with infinitesimal automorphisms of weak Carrollian manifolds where the Lie algebra is infinite-dimensional (see Duval et al. [9]).
@@
A super-null Riemannian manifold $(M,g,Q,P)$ such that $P$ is Killing, i.e., $\mathcal{L}_{P}g=0$, is referred to as a static super-null Riemannian manifold.
Note that for any static super-null Riemannian manifold $P\in\mathfrak{snull}_{0}(M,g,Q,P)$ as $\mathcal{L}_{P}Q=[P,Q]=0$. Thus, the super-null Lie algebra of a static super-null Riemannian manifold consists of at least $P$ and $D$. As standard, we have a representation of the $N=1$, $d=1$ supertranslation algebra given by [D,D]=2P[D,D]=-2P and $[P,Q]=0$. Thus, we have established the following.
@@
Let $(M,g,Q,P)$ be a static super-null Riemannian manifold, then the $N=1$, $d=1$ supertranslation algebra is a Lie subsuperalgebra of $\mathfrak{snull}_{(}M,g,Q,P)$ realised by the vector fields $P$ and $D$.
@@
i.e., we have the usual Minkowski spacetime of signature $(3,1)$. Clearly, $P$ is Killing and we have a static super-null Riemannian manifold, so the super-null Lie algebra contains $D$. The isometries of $g_{red}$ are given by the Poincaré Lie algebra, i.e., $\mathfrak{iso}(g_{red})\simeq\mathfrak{so}(3,1)\ltimes{\mathbb{R}}^{4}$. However, we require the Lie subalgebra generated by $P_{red}=\partial_{t}$ and the other generators of the Poincaré Lie algebra that commute with $P_{red}$. The remaining transformations are spatial rotations, spatial translations, and temporal translations. Importantly, there are no boosts. Thus,
@@
Let $(M_{0},g_{0})$ be a (pseudo-)Riemannian manifold with $\dim M_{0}\geq 2$, and let $\mathfrak{iso}(g_{0})$ be its isometry Lie algebra. We then equip $M_{red}:=M_{0}\times{\mathbb{R}}$ with the product metric $g_{red}:=g_{0}\oplus g_{\mathbb{R}}$, where $g_{\mathbb{R}}$ is the standard constant metric on $\mathbb{R}$}. An established result is that $\mathfrak{iso}(g_{0}\oplus g_{\mathbb{R}})\simeq\mathfrak{iso}(g_{0})\oplus\mathfrak{iso}(g_{\mathbb{R}})$. Recall that $\mathfrak{iso}(g_{R})$ consists of just translations and thus is identified with $\mathfrak{u}(1)$. Then $M:=M_{0}\times{\mathbb{R}}^{1|1}$ is a super-null Riemannian manifold whose degenerate metric written in Shander coordinates is
@@
$\mathfrak{snull}(M_{0}\times{\mathbb{R}}^{1|1})\simeq\big(\mathfrak{iso}(M_{0})\oplus\mathfrak{u}(1)\big)\oplus_{\textrm{E}xt}{\mathbb{R}}^{0|1}\,$,
@@
2.4. Connections on Super-Null Riemannian Manifolds
@@
Let $(M,g,Q,P)$ be a super-null Riemannian manifold. An affine connection on the supermanifold $M$ is said to be
@@
Let $(M,g,Q,P)$ be a super-null Riemannian manifold equipped with an affine connection $\nabla$. If the affine connection is supersymmetry compatible, then $\nabla$ cannot be torsionless.
@@
Let $(M,g,Q,P)$ be a super-null Riemannian manifold equipped with an affine connection $\nabla$. If the affine connection is supersymmetry compatible, then $Q\in\ker\big(R_{\nabla}(X,Y)\big)$ for all pairs of vector fields $X,Y\in{Vect}(M)$.
@@
Supersymmetry compatible connections exist on any super-null Riemannian manifold.
@@
Let $(M,g,Q,P)$ be a super-null Riemannian manifold equipped with a metric compatible connection. For every $X\in{Vect}(M)$ there exists a function $f_{X}\in C^{\infty}(M)$ ($\widetilde{f_{X}}=\widetilde{X}$), such that $\nabla_{X}Q=f_{X}\,Q$.
@@
Let $(M,g,Q,P)$ be a super-null Riemannian manifold equipped with an affine connection $\nabla$. If the affine connection is metric compatible, then $\nabla$ cannot be torsionless.
@@
Let $(M,g,Q,P)$ be a super-null Riemannian manifold equipped with an affine connection. If the affine connection is metric compatible, then $Q$ is a generalised eigenvector of $R_{\nabla}(X,Y)$ for all pairs of vector fields $X,Y\in{Vect}(M)$.
@@
does not fully determine a metric compatible connection. While $\langle\nabla_{X}Y|Z\rangle$ is fully determined for a given metric compatible connection (assuming at least one exists), $\nabla_{X}Y$ is only determined up to a vector in $\ker(g)$. Thus, we cannot expect a direct analogue of the fundamental theorem of Riemannian supergeometry for super-null Riemannian manifolds. Nonetheless, we have the following theorem.
@@
Metric compatible affine connections exist on any super-null Riemannian manifold.
@@
Compatible affine connections exist on any super-null Riemannian manifold.
@@ 3. Concluding Remarks
An odd analogue of a Carrollian manifold has been constructed and studied; we refer to such supermanifolds as super-null manifolds. It is important to note that underlying a super-null manifold is a pseudo-Riemannian manifold and not a Carrollian manifold. We have shown that supersymmetry compatible and metric compatible connections exist on any super-null Riemannian manifold, and importantly, that compatible connections always exist, i.e., affine connections that are both supersymmetry compatible and metric compatible can always be constructed. It is the freedom in defining a connection, thanks to the degeneracy of the metric, that allows these two conditions to be simultaneously satisfied. It was argued by Bekaert & Morand [2] that only invariant Carrollian manifolds, i.e., Carrollian manifolds for which the fundamental/Carrollian vector field is Killing, can admit torsion-free compatible connections. This is in stark contrast with super-null Riemannian manifolds, where the supersymmetry generator $Q$ cannot be Killing, and compatible connections must carry torsion. The two geometries are fundamentally different; the non-integrability of the kernel of the super-case is the root of these differences.
The Lie superalgebra of infinitesimal automorphisms of a super-null Riemannian manifold, which we referred to as the super-null Lie algebra, has been studied. An interesting result is that this Lie superalgebra is finite-dimensional and tightly tied to the Lie algebra of Killing vector fields of the reduced metric. In the classical setting of weak Carrollian manifolds, the Lie algebra of infinitesimal automorphisms is infinite-dimensional. When the extra condition of preserving a compatible connection is imposed, the Lie algebra becomes finite-dimensional; see Duval et al. [9] and references therein for details.
While the main motivation for this work is rooted in mathematical curiosity, further explicit examples of super-null Riemannian manifolds are desirable and could expose applications thereof in physics. It may be possible to formulate a superspace version of non-holonomic supermechanics where the dynamics generated by $Q$ could depend on the extra even coordinates understood as “external parameters”; which might be interpreted as the classical background fields or slowly varying time-dependent parameters of the physical system. This could provide a novel way to study supersymmetric mechanical systems with evolving external parameters. For example, there may be supersymmetric analogues of geometric phases, such as the Hannay angle (see [18]), underlying these models.